OpenStudy (anonymous):

how far can a car travel in 40 mins if it is traveling at an average speed of 50 miles per hour

OpenStudy (anonymous):

D= r t r = rate t = time

OpenStudy (anonymous):

40 mins = 2/3 hours

OpenStudy (anonymous):

50 * 2/3 = 100/3 miles

OpenStudy (anonymous):

a.2000mi b.1.25mi c.33.3mi d.0.8mi

OpenStudy (anonymous):

c